Research Proposal Form

Date: June, 1997

1 Contractor or Principal Investigator

• Organization

• Address of the Organization

• Title

• Name

• Contact Address TEL FAX e-mail

2 Title of the Proposed Project

(1) Title of the Proposed Project:

(2) Research Period (one-year project or two-year project):

(3) Background of the research (about a half page):

(4) Purpose of the research (about a half page):

(5) Contents of the research (around 1 to 2 pages):

3 Project Organization/Research

Method/Scedule

(1) Project Organization

Name Affiliation e-mail address

Principal Investigator Corporate Researcher

(2) Research Method

(3) Research Schedule: schedule for two years for two-year project

4 Resulting Software

(1) The name of the software

(2) Functions and features of the software (more than a half page)

(3) Organization of the software (about a half page)

• Organization of the software

• If the resulting software will be a part of larger system, please explain the relation of the software to the total system.

• If the resulting software will be improvement of existing software by adding new functions and etc., please explain what kind of functions are added.

(4) If the software is related to specific IFS programs, please specify their names.

(5) Please specify the programming language/OS/software packages for the execution of the software, and please explain the portability of resulting software.

(6) Expected size of the software (restricted to newly developed part)

(7) Expected way of use of the software (more than a half page) • How this software will be used, who will be expected users.

• Advantage of the software from users’ point of view.

(8) Expected intermediate result at the end of the first fiscal year (if the project is a two-year project)

(9) Documents which will be added to the software

• Specification, User manual, etc.

5 Bases of Funding Money

Item  Amount Bases Remarks

Personnel Expenses Traveling Expenses

Materials Purchase Expenses Expendables Expenses Miscellaneous Expences